# Nuclear-physics implications of a subsolar neutron star gravitational-wave observation

### v1

@@ Abstract
Detecting a compact subsolar object would have profound implications in physics, the reach of which depends on the nature of the object. Here we explore such consequences for a putative subsolar-mass gravitational wave event detected by the LIGO-Virgo-KAGRA Collaboration. We forecast that the nature of a subsolar binary (made of light neutron stars, primordial black holes, or more exotic compact objects) can be inferred with a great statistical confidence level already during the ongoing fourth observing run, based on the large tidal deformability effects on the signal. The detection of a primordial black hole would have implications for cosmology and dark matter scenarios, while the measurement of the tidal deformability of a subsolar neutron star could rule out or confirm the existence of strange stars made of quarks.
@@
Introduction. Since the groundbreaking detection of a gravitational wave (GW) event by LIGO in 2015 [1], the LIGO-Virgo-KAGRA (LVK) Collaboration has reported nearly a hundred GW detections originating from the coalescence of compact binary systems [2, 3, 4]. Current GW interferometers are capable of identifying compact binary coalescences in a wide range of masses and are sensitive to mergers of hypothetical subsolar mass (SSM) compact objects (with mass $m<M_{\odot}$). Conventional formation models [5] suggest that astrophysical compact objects, such as black holes (BHs) and neutron stars (NSs), should have masses exceeding the solar mass (see, e.g., [6, 7, 8, 9] for recent simulations), although white dwarfs and NSs can theoretically be subsolar [10, 11, 12, 13].
@@
The detection of SSM compact objects could either indicate a novel formation mechanism beyond the standard-model stellar core-collapse scenarios [13], or imply evidence of new physics such as primordial black holes (PBHs) [14, 15, 16, 17] (see e.g. [18, 19] for recent reviews) or more exotic objects [20] such as Q-balls/boson stars (BSs) [21, 22, 23], fermion-soliton stars [24, 25, 26], etc. Whatever the case, such a discovery would have important implications for astrophysics, cosmology, and fundamental physics.
@@
Observing SSM objects in compact binary coalescing systems is therefore a primary target for current and future [27] GW interferometers. Several GW searches for compact coalescing binaries with at least one SSM component have been performed using LVK data, but no conclusive detections have been made [28, 29, 30, 31]. However, in the last concluded LVK observing run, O3b, three candidates of SSM binary BH events were identified [31], and one candidate from the O2 data was re-examined in [32]. None of these candidates were confirmed due to their relatively high false alarm rates.
@@
Figure 1: The posterior distribution ($50\%$ and $90\%$ C.L.) of mass and tidal deformability in a simulated observation of a SSM binary NS during O4 (O5) with ${\rm SNR}=25$ (${\rm SNR}=44$), assuming strange quark matter EoS. Different curves correspond to theoretical predictions for different NS EoS and for BSs. As discussed below, such a measurement and Bayesian EoS selection [33] during O4/O5 would exclude the ordinary EoS with normal $npe\mu$ matter, and would also exclude the PBH and BS origin of the binary.
@@
The scope of this work is to explore the following question: What would be the implications of a putative SSM GW event? As we will show, already the ongoing LVK O4 run (and even more so the next O5 run scheduled to start in 2026) has sufficient sensitivity to discern among various scenarios, all with very impactful consequences.
@@
In particular, we present the results of thorough Bayesian inferences on the candidate event SSM200308 and on projected O4 and O5 data, using several models for SSM objects, including PBHs, light NSs, and BSs, the latter taken as a representative model of exotic compact objects. We show that already SSM200308, if interpreted as a real event, would decisively exclude some models of NSs and BSs compared to the PBH hypothesis, based on the large tidal deformability effects on the waveform [34, 35]. We will show that in O4 this distinction would be crystal-clear, confirming or ruling out a putative PBH event. In the latter case, using O4-O5 data it would be possible to use tidal-deformability measurements to decisively confirm/exclude light NSs with strange quark matter [36, 37], giving a decisive contribution to the quest for the equation of state (EoS) of nuclear matter at ultrahigh densities (see, e.g., [38, 39, 40]). As anticipated in Fig. 1 and discussed below, this compelling opportunity is provided by the fact that, in the SSM range, quark stars have a much smaller radius (and hence much smaller tidal deformability) than ordinary-EoS NSs, and that overall the tidal deformability for SSM NSs is very large [34], so can be more easily measured than for heavier NSs.
@@
Methodology & Models. We use the same methodology implemented in Ref. [34], summarized below and detailed in the Supplemental Material (SM). In particular, the Bayesian inference on both real and synthetic GW data is done with the public software BILBY [41]. We consider quasicircular binaries, for which the waveform parameters common in all models are
@@
[…]
We consider 6 different models:
@@
i) BH1: the same IMRPhenomPv2 waveform [42] used in [43] to analyze SSM200308. This model assumes a BH binary and includes spin precession, so it augments the parameters listed above with the spin angles. We adopt the same priors of [43], with extended ranges in some cases, see the SM for more details.
@@
ii) BH2: this is a simplified BH binary model based on the TaylorF2 waveform [44], including up to 3.5 post-Newtonian (PN) terms [45, 46, 47, 48] in the point-particle phase, and up to 4PN spin effects (see Refs. [49, 50] and references therein for details, e.g. [51, 52, 53, 54, 55, 56, 57, 58]) assuming spins aligned with the orbital angular momentum.
@@
iii) Agnostic: TaylorF2 waveform augmented with the 5PN and 6PN tidal terms in the phase [59], parametrized in terms of the dimensionless tidal deformabilities $\Lambda_{i}$ ($i=1,2$) of the binary components [34], and with arbitrary spin-induced quadrupole terms at 2PN [54], parametrized in terms of normalized quadrupole moments $k_{i}$ (with $k_{i}=1$ in the BH case). Finally, the model includes the tapering function introduced in [34] (see also [60]) to account for tidal disruption of objects composing the binary, which is parametrized in terms of a cutoff frequency $f_{\text{\tiny cutoff}}=\tilde{\lambda}_{f}f_{\text{\tiny ISCO}}$, where $f_{\text{\tiny ISCO}}=4.4\,{\rm kHz}[M_{\odot}/(m_{1}+m_{2})]$ is the frequency of the innermost stable circular orbit of the binary, and a slope $f_{\text{\tiny slope}}=\delta\tilde{\lambda}_{f}f_{\text{\tiny ISCO}}$, which controls the smoothness of the tapering, fixed to $\delta\tilde{\lambda}_{f}=10^{-2}$. In practice, this model augments BH2 with 5 extra parameters: $\Lambda_{i}$, $k_{i}$, $\tilde{\lambda}_{f}$}. As usually done, we perform the inference in terms of the effective parameters $\tilde{\Lambda}(\Lambda_{i},m_{i})$ and $\delta\tilde{\Lambda}(\Lambda_{i},m_{i})$, associated to the 5PN and 6PN order contribution to the GW phase (see e.g. [34] for their explicit expressions). Being the model agnostic, we implement large uniform priors on the extra parameters (in particular $\Lambda_{i}$, $k_{i}$ can be either positive or negative) and the same priors used in BH2 for the standard parameters.
@@
iv) NS1: this model is based on Agnostic but enforces some relations among parameters, motivated by concrete NS models. In particular, we implement the approximately EoS-independent “Love-Q” relations [61, 62, 63, 64] which relate $k_{i}$ to $\Lambda_{i}$. Finally, we assume a cutoff frequency given by the Roche overflow of the secondary star [65, 66, 67, 68]. Fitting the numerical simulations presented in Ref. [69], we estimate [34]
@@
v) NS2: similar to NS1, but we ignore the Love-Q relations and assume the same spin-induced quadrupole moments of a BH ($k_{i}=1$).
vi) BS: this model implements a system of binary BSs with strong quartic interactions in the potential, for which the template was already developed in Ref. [70]. In this case the parameters ${\bm{\theta}}$ are supplemented by $M_{B}$, which is the mass coupling constant related to the quartic potential [71, 70]. We also insert the waveform smoothing, which is controlled by the Roche frequency [70]:
@@
Analysis of SSM200308. With a false-alarm rate of 1 per 5 years, SSM200308 stands out as the most significant SSM candidate identified by the GstLAL pipeline [73] in both LIGO Hanford and LIGO Livingston detectors. Although SSM200308 did not generate a trigger in Virgo with a signal-to-noise ratio (SNR) above the single detector threshold, Virgo was operational at that time, and its data is included in our analysis.
@@
SSM200308 was initially analyzed under the assumption that it resulted from the merger of binary BHs, with reported masses of $m_{1}=0.62^{+0.46}_{-0.20}\,M_{\odot}$ and $m_{2}=0.27^{+0.12}_{-0.10}\,M_{\odot}$ at a redshift of $z=0.02^{+0.01}_{-0.01}$ (90% C.I.) [43]. Recent forecasts, however, showed that the ongoing observing run O4 could potentially distinguish, through tidal deformability measurements, whether the trigger originated from the merger of two PBHs or other SSM candidates [34]. Motivated by these findings, we conduct Bayesian inferences on the SSM200308 data under the different hypothesis presented above. Our results do not alter the low significance of SSM200308 and we do not claim evidence of this being an actual event, as it may also have originated from environmental or instrumental noise. Rather, we show that such long-duration and low-mass signals could give valuable insight to distinguish different hypothesis, even more so with the anticipated improvements in sensitivity of future observing runs.
@@
The results of our analysis are summarized in Table 1, where we show the $90\%$ C.I. for the most representative parameters. The posterior distributions of the masses are compatible in all models and at least the secondary mass $m_{2}$ could always be statistically inferred to be subsolar and to satisfy the condition $m_{2}+3\Delta m_{2}<M_{\odot}$ [49], where $\Delta m_{2}$ is its standard deviation. The results with BH1 are compatible with the ones reported in Ref. [43]. We adopt this as the baseline to compare the Bayes factors between different models. For BH1, we find that the data prefer a positive effective spin $\chi_{\text{\rm eff}}$ with high significance, while the precession parameter $\chi_{\rm p}$ is only weakly constrained. Assuming the BH2 model, which neglects spin precession, we find similar results to the BH1 case. Therefore, given also the Bayes factors reported in the table, we conclude there is limited evidence for precession in current data.
@@
With the Agnostic hypothesis, we observe that the posteriors of the dimensionless quadrupoles $k_{i}$ are broad, which is compensated by a low $\chi_{\rm eff}$, so that the corresponding 2PN phase term, controlled by the combinations $k_{i}\chi_{i}^{2}$, remains small. Also, the tidal deformabilities are large enough to include all possible origins of the SSM candidate. This can be due to the limited sensitivity at large frequencies during the O3 run, as well as the strong preference for a cut-off frequency below the ISCO, as indicated by the the tapering parameter $\log_{10}\tilde{\lambda}_{f}$ being inferred to be smaller then zero with large significance. While the latter would disfavor the binary BH interpretation, the relative $\log_{10}$ Bayes factor to the BH1 case is negative, possibly due to weakly informative data and the larger number of parameters in the Agnostic model.
@@
For what concerns the models NS1 and NS2, while the latter has a Bayes factor similar to BH1, the former shows a decisive negative difference. We expect that the inclusion of the Love-Q relations is responsible for generating instabilities at high spins. Note that, despite NS2 not being statistically disfavored, it infers highly spinning binary components which could be in tension with a NS interpretation, since old NSs in the late stages of the inspiral are expected to rotate slowly [74].
@@
Finally, under the BS hypothesis one would place the coupling $M_{B}$ close to $10M_{\odot}$, but this scenario is decisively disfavored by the data compared to other models, probably for the same reason of NS1 as in the BS model we implement its own Love-Q relations [75, 70].
@@
In Table 2, we show the corresponding Bayes factors. We observe that O4 sensitivity is sufficient to distinguish a PBH signal from the NS one with similar masses, especially at higher SNR. Our results strongly suggest that, should a PBH SSM binary be detected already in O4, it can be confidently distinguished from a SSM NS binary, and vice versa. Similar conclusions apply to the BS model (not shown), which can be decisively excluded in the case of a BH or NS injection. Furthermore, as shown in Fig. 2, interpreting a SSM BH binary signal with the wrong NS model (or vice versa) would introduce a significant bias in both masses and tidal deformabilities, with the posteriors of the latter effectively being compatible with zero (with the C.L. cut to finite small values only due to the assumed flat prior on $\Lambda_{i}$, see discussion below). This shows that a Bayesian model selection for these sources is crucial to properly identify the nature of the objects.
Detector BH2 $\rightarrow$ NS2 NS2 $\rightarrow$ BH2
O4, SNR=15 -3.1 -2.9
O4, SNR=25 -5.1 -11.0
Table 2: $\log_{10}$ Bayes factors to assess the PBH or NS origin of a putative SSM event, normalized to the case where the injection and recovery assume the same model (i.e. negative values disfavor the wrong hypothesis).
Cosmology & Nuclear Physics implications. Having assessed that the nature of a SSM can be confidently identified, we now turn our attention to the implications of such putative detection.
In the case of a SSM BH binary detected during O4, our results show that one can rule out the NS interpretation with decisive Bayesian evidence, mostly through the observation of the GW signal beyond the expected disruption frequency (i.e. $\tilde{\lambda}_{f}\sim 1$). [^1] We are neglecting here the possible impact of environmental effects which could generate non-zero tidal effects in the relatively low-frequency range of the inspiral signal [76, 77, 60, 78, 79, 80, 81]. One could then infer the corresponding PBH abundance $f_{\rm PBH}$ (defined as the fraction of dark matter energy density made of PBHs) which controls the merger rate of similar SSM events. Following the analysis of Ref. [82], based on a model saturating the GWTC-3 upper bound, one would need at least $f_{\rm PBH}\gtrsim{\cal O}(10^{-2})$ to explain such a SSM event, close but not excluded by the most recent bounds from missing microlensing signatures in OGLE datasets [83]. Due to the finite width of the PBH mass distribution (at least dictated by the features of the critical collapse), one could then expect PBH events also to contaminate the stellar mass range. While such observation would point towards an important role of PBHs in the evolution of large-scale structures (see e.g. [84, 85, 86, 87, 88]), it would also strengthen bounds from indirect detection of particle dark matter, which is largely incompatible with an abundant population of compact objects boosting the annihilation J-factors [89, 90, 91, 92, 93]. Finally, the existence of a PBH population would require specific setups in the early universe, either deviating from the vanilla slow roll inflationary scenario [94] or other formation mechanisms based on new physics (see e.g. Ref. [95]). Another nonprimordial explanation could be subsolar BHs formed from stellar transmutation triggered by dark matter accretion [96, 97, 98, 99, 100, 101, 102], which anyway require new physics.

### v2

# Cosmology and nuclear-physics implications of a subsolar gravitational-wave event

@@ Abstract
Detecting a compact subsolar object would have profound implications in physics, the reach of which depends on the nature of the object. Here, we explore such consequences for a putative subsolar-mass gravitational wave event detected by the LIGO-Virgo-KAGRA Collaboration. We forecast that the nature of a subsolar binary (made of light neutron stars, primordial black holes, or more exotic compact objects) can be inferred with a great statistical confidence level already during the ongoing fourth observing run, based on the large tidal deformability effects on the signal. The detection of a primordial black hole would have implications for cosmology and dark matter scenarios, while the measurement of the tidal deformability of a subsolar neutron star could rule out or confirm the existence of strange stars made of quarks.
@@ I Introduction
Since the groundbreaking detection of a gravitational wave (GW) event by LIGO in 2015 [1], the LIGO-Virgo-KAGRA (LVK) Collaboration has reported nearly a hundred GW detections originating from the coalescence of compact binary systems [2, 3, 4]. Current GW interferometers are capable of identifying compact binary coalescences in a wide range of masses and are sensitive to mergers of hypothetical subsolar mass (SSM) compact objects (with mass $m<M_{\odot}$). Conventional formation models [5] suggest that astrophysical compact objects, such as black holes (BHs) and neutron stars (NSs), should have masses exceeding the solar mass (see, e.g., Refs. [6, 7, 8, 9, 10] for recent simulations), although white dwarfs and NSs can theoretically be subsolar [11, 12, 13, 14].
@@
The detection of SSM compact objects could either indicate a novel formation mechanism beyond the standard-model stellar core-collapse scenarios [14, 10] or imply evidence of new physics such as primordial black holes (PBHs) [15, 16, 17, 18] (see, e.g., Refs. [19, 20] for recent reviews) or more exotic objects [21] such as Q-balls/boson stars (BSs) [22, 23, 24], fermion-soliton stars [25, 26, 27], etc. Whatever the case, such a discovery would have important implications for astrophysics, cosmology, and fundamental physics.
Examples of SSM compact objects may have already been observed. HESS J1731–347 [28] is a candidate NS with mass $0.77^{+0.20}_{-0.17}M_{\odot}$, while the candidate object reported in Ref. [29] was possibly identified as a white dwarf with mass $0.20^{+0.01}_{-0.01}M_{\odot}$.
@@
Observing SSM objects in compact binary coalescing systems is, therefore a primary target for current and future [30] GW interferometers. Several GW searches for compact coalescing binaries with at least one SSM component have been performed using LVK data, but no conclusive detections have been made [31, 32, 33, 34]. However, in the last concluded LVK observing run, O3b, three candidates of SSM binary BH events were identified [34], and one candidate from the O2 data was reexamined in Ref. [35]. None of these candidates was confirmed due to their relatively high false alarm rates.
@@
Figure 1: The posterior distribution ($50\%$ and $90\%$ C.L.) of mass and tidal deformability in a simulated observation of a SSM binary NS during O4 (O5) with ${\rm SNR}=25$ (${\rm SNR}=44$), assuming strange quark matter EoS SQM3. Different curves correspond to theoretical predictions for different NS EoS and for BSs. As discussed below, such a measurement and Bayesian EoS selection [36] during O4/O5 would exclude the ordinary EoS with normal $npe\mu$ matter and would also exclude the PBH and BS origin of the binary.
@@
The scope of this work is to explore the following question: What would be the implications of a putative SSM GW event? As we will show, the ongoing LVK O4 run (and even more so the next O5 run scheduled to start in 2026) has sufficient sensitivity to discern among various scenarios, all with very impactful consequences.
@@
In particular, we present the results of thorough Bayesian inferences on the candidate event SSM200308 and on projected O4 and O5 data, using several models for SSM objects, including PBHs, light NSs, and BSs, the latter taken as a representative model of exotic compact objects. We show that already SSM200308, if interpreted as a real event, would decisively exclude some models of NSs and BSs compared to the PBH hypothesis, based on the large tidal deformability effects on the waveform [37, 38]. We will show that in O4 this distinction would be crystal clear, confirming or ruling out a putative PBH event. In the latter case, using O4–O5 data it would be possible to use tidal-deformability measurements to decisively confirm/exclude light NSs with strange quark matter [39, 40], giving a decisive contribution to the quest for the equation of state (EoS) of nuclear matter at ultrahigh densities (see, e.g., Refs. [41, 42, 43]). As anticipated in Fig. 1 and discussed below, this compelling opportunity is provided by the fact that, in the SSM range, quark stars have a much smaller radius (and hence much smaller tidal deformability) than ordinary-EoS NSs, and that overall the tidal deformability for SSM NSs is very large [37] and so can be more easily measured than for heavier NSs.
II Methodology and Models
@@
We use the same methodology implemented in Ref. [37], summarized below and detailed in Appendix A. In particular, the Bayesian inference on both real and synthetic GW data is done with the public software BILBY [44]. We consider quasicircular binaries, for which the waveform parameters common in all models are
@@
Table 1: 90% C.I. inferred for the trigger event SSM200308 using different waveform models: BH1 (no tidal effects, precession), BH2 (no tidal effects, no precession), Agnostic (generic tidal deformability, quadrupoles, and tidal cutoff frequency), NS1/2 (two waveform models for light binary NS), and BS (a model for BS binaries). The Bayes factors are defined as $\mathcal{B}_{i}={\cal Z}_{i}/{\cal Z}_{{\rm BH1}}$, where ${\cal Z}_{i}$ is the evidence of a given model and $i=\{{\rm BH2,Agnostic,NS1,NS2,BS}\}$. According to Jeffreys’s scale criterion [45], a $\log_{10}$ Bayes factors greater than $1$ (respectively, $2$) would imply a strong (respectively, decisive) Bayesian evidence in favor of a given model relative to BH1.
We consider six different models:
@@
i) BH1. The same IMRPhenomPv2 waveform [46] used in Ref. [47] to analyze SSM200308. This model assumes a BH binary and includes spin precession, so it augments the parameters listed above with the spin angles. We adopt the same priors of Ref. [47], with extended ranges in some cases; see Appendix A for more details.
@@
ii) BH2. This is a simplified BH binary model based on the TaylorF2 waveform [48], including up to 3.5 post-Newtonian (PN) terms [49, 50, 51, 52] in the point-particle phase and up to 4PN spin effects (see Refs. [53, 54] and references therein for details, e.g., Refs. [55, 56, 57, 58, 59, 60, 61, 62]) assuming spins aligned with the orbital angular momentum.
@@
iii) Agnostic. TaylorF2 waveform augmented with the 5PN and 6PN tidal terms in the phase [63], parametrized in terms of the dimensionless tidal deformabilities $\Lambda_{i}$ ($i=1,2$) of the binary components [37], and with arbitrary spin-induced quadrupole terms at 2PN [58], parametrized in terms of normalized quadrupole moments $k_{i}$ (with $k_{i}=1$ in the BH case). As discussed in Ref. [37], SSM binaries have high merger frequencies outside the bandwidth of current detectors. The observed signal originates from the early inspiral phase, where the TaylorF2 approximant remains accurate. This approach also avoids the need for more sophisticated binary NS waveform models, which are inaccurate in the SSM range [64]. Finally, our model includes the tapering function introduced in Ref. [37] (see also Ref. [65]) to account for tidal disruption of objects composing the binary, which is parametrized in terms of a cutoff frequency $f_{\text{\tiny cutoff}}=\tilde{\lambda}_{f}f_{\text{\tiny ISCO}}$, where $f_{\text{\tiny ISCO}}=2.2\,{\rm kHz}[M_{\odot}/(m_{1}+m_{2})]$ is the frequency of the innermost stable circular orbit of the binary, and a slope $f_{\text{\tiny slope}}=\delta\tilde{\lambda}_{f}f_{\text{\tiny ISCO}}$, which controls the smoothness of the tapering, fixed to $\delta\tilde{\lambda}_{f}=10^{-2}$. In practice, this model augments BH2 with five extra parameters: $\Lambda_{i}$, $k_{i}$, $\tilde{\lambda}_{f}$}. As usually done, we perform the inference in terms of the effective parameters $\tilde{\Lambda}(\Lambda_{i},m_{i})$ and $\delta\tilde{\Lambda}(\Lambda_{i},m_{i})$, associated to the 5PN and 6PN order contribution to the GW phase (see, e.g., Ref. [37] for their explicit expressions). With the model being agnostic, we implement large uniform priors on the extra parameters (in particular $\Lambda_{i}$ and $k_{i}$ can be either positive or negative) and the same priors used in BH2 for the standard parameters.
@@
iv) NS1. This model is based on Agnostic but enforces some relations among parameters, motivated by concrete NS models. In particular, we implement the approximately EoS-independent “Love-Q” relations [66, 67, 68, 69] which relate $k_{i}$ to $\Lambda_{i}$. Finally, we assume a cutoff frequency given by the Roche overflow of the secondary star [70, 71, 72, 73]. Fitting the numerical simulations presented in Ref. [74], we estimate [37]
@@
v) NS2. Similar to NS1, but we ignore the Love-Q relations and assume the same spin-induced quadrupole moments of a BH ($k_{i}=1$).
vi) BS. This model implements a system of binary BSs with strong quartic interactions in the potential, for which the template was already developed in Ref. [75]. In this case, the parameters ${\bm{\theta}}$ are supplemented by $M_{B}$, which is the mass coupling constant related to the quartic potential [76, 75]. We also insert the waveform smoothing, which is controlled by the Roche frequency [75],
@@ III Analysis of SSM200308
With a false-alarm rate of 1 per 5 years, SSM200308 stands out as the most significant SSM candidate identified by the GstLAL pipeline [78] in both LIGO Hanford and LIGO Livingston detectors. Although SSM200308 did not generate a trigger in Virgo with a signal-to-noise ratio (SNR) above the single detector threshold, Virgo was operational at that time, and its data are included in our analysis.
@@
SSM200308 was initially analyzed under the assumption that it resulted from the merger of binary BHs, with reported masses of $m_{1}=0.62^{+0.46}_{-0.20}\,M_{\odot}$ and $m_{2}=0.27^{+0.12}_{-0.10}\,M_{\odot}$ at a redshift of $z=0.02^{+0.01}_{-0.01}$ (90% C.I.) [47]. Recent forecasts, however, showed that the ongoing observing run O4 could potentially distinguish, through tidal deformability measurements, whether the trigger originated from the merger of two PBHs or other SSM candidates [37]. Motivated by these findings, we conduct Bayesian inferences on the SSM200308 data under the different hypotheses presented above. Our results do not alter the low significance of SSM200308, and we do not claim evidence of this being an actual event, as it may also have originated from environmental or instrumental noise. Rather, we show that such long-duration and low-mass signals could give valuable insight to distinguish different hypothesis, even more so with the anticipated improvements in sensitivity of future observing runs.
@@
The results of our analysis are summarized in Table 1, where we show the $90\%$ C.I. for the most representative parameters. The posterior distributions of the masses are compatible in all models and at least the secondary mass $m_{2}$ could always be statistically inferred to be subsolar and to satisfy the condition $m_{2}+3\Delta m_{2}<M_{\odot}$ [53], where $\Delta m_{2}$ is its standard deviation. The results with BH1 are compatible with the ones reported in Ref. [47]. We adopt this as the baseline to compare the Bayes factors between different models. For BH1, we find that the data prefer a positive effective spin $\chi_{\text{\rm eff}}$ with high significance, while the precession parameter $\chi_{\rm p}$ is only weakly constrained. Assuming the BH2 model, which neglects spin precession, we find similar results to the BH1 case. Therefore, given also the small Bayes factor in the table ($\log_{10}{\cal B}=0.31$), we conclude there is limited evidence for precession in current data.
@@
With the Agnostic hypothesis, we observe that the posteriors of the dimensionless quadrupoles $k_{i}$ are broad, which is compensated by a low $\chi_{\rm eff}$, so that the corresponding 2PN phase term, controlled by the combinations $k_{i}\chi_{i}^{2}$, remains small. Also, the tidal deformabilities are large enough to include all possible origins of the SSM candidate. This can be due to the limited sensitivity at large frequencies during the O3 run as well as the strong preference for a cutoff frequency below the ISCO, as indicated by the tapering parameter $\log_{10}\tilde{\lambda}_{f}$ being inferred to be smaller than zero with large significance. While the latter would disfavor the binary BH interpretation, the relative $\log_{10}$ Bayes factor to the BH1 case is negative, possibly due to weakly informative data and the larger number of parameters in the Agnostic model.
@@
For what concerns the models NS1 and NS2, while the latter has a Bayes factor similar to BH1, the former shows a decisive negative difference according to Jeffreys’s scale [45]. We expect that the inclusion of the Love-Q relations is responsible for generating instabilities at high spins. Note that, despite NS2 not being statistically disfavored, it infers highly spinning binary components which could be in tension with a NS interpretation, since old NSs in the late stages of the inspiral are expected to rotate slowly [79].
@@
Finally, under the BS hypothesis, one would place the coupling $M_{B}$ close to $10M_{\odot}$, but also this scenario is decisively disfavored by the data compared to other models, probably for the same reason of NS1 as in the BS model we implement its own Love-Q relations [80, 75].
IV Forecasts for LVK O4/O5
@@
In Table 2, we show the corresponding Bayes factors, focusing on two models (BH2 and NS2), which could not be discriminated according to the Bayesian analysis shown in Table 1. We observe that O4 sensitivity is sufficient to distinguish a PBH signal from the NS one with similar masses, especially at higher SNR. Our results strongly suggest that, should a PBH SSM binary be detected already in O4, it can be confidently distinguished from an SSM NS binary and vice versa. Because the two cases are already distinguishable in O4, O5 will be even more capable of discriminating between them. As an estimate, a SNR of 15 in O4 corresponds roughly to an SNR of 25 in O5 [37]. Similar conclusions apply to the BS model (not shown), which can be even more strongly excluded in the case of a BH or NS injection. Furthermore, as shown in Fig. 2, interpreting a SSM BH binary signal with the wrong NS model (or vice versa) would introduce a significant bias in both masses and tidal deformabilities. This shows that performing model selection for these sources is crucial to properly identifying the nature of the objects.
Detector $BH2$rightarrow $NS2$ $NS2$rightarrow $BH2$
SNR=15 -3.1 -2.9
SNR=25 -5.1 -11.0
Table 2: $\log_{10}$ Bayes factors to assess the PBH or NS origin of a putative SSM event, normalized to the case where the injection and recovery assume the same model (i.e., negative values disfavor the wrong hypothesis). We assumed O4 sensitivity, and results for O5 at the same SNR would be very similar.
@@
V Cosmology and Nuclear Physics implications
Having assessed that the nature of a SSM can be confidently identified, we now turn our attention to the implications of such putative detection.
V.1 Primordial hypothesis
In the case of a SSM BH binary detected during O4, our results show that one can rule out the NS interpretation with decisive Bayesian evidence, mostly through the observation of the GW signal beyond the expected disruption frequency (i.e. $\tilde{\lambda}_{f}\sim 1$). [^1] We are neglecting here the possible impact of environmental effects which could generate nonzero tidal effects in the relatively low-frequency range of the inspiral signal [81, 82, 65, 83, 84, 85, 86]. These would yield frequency-dependent features that can be disentangled from those of NS tidal deformabilities. One could then infer the corresponding PBH abundance $f_{\rm PBH}$ (defined as the fraction of dark matter energy density made of PBHs) that controls the merger rate of similar SSM events. Following the analysis of Ref. [87], based on a model saturating the GWTC-3 upper bound, one would need an abundance of order $f_{\rm PBH}\sim{\cal O}(10^{-2})$ to explain such a SSM event, a value that depends on the possible assumed mass distribution. This is close but not excluded by the most stringent bounds in this mass range arising from missing microlensing signatures in OGLE datasets [88]. Because of the finite width of the PBH mass distribution required at least by the features of the critical collapse, one could then expect PBH events also to contaminate the stellar mass range. While such observation would point toward an important role of PBHs in the evolution of large-scale structures (see, e.g., Refs. [89, 90, 91, 92, 93]), it would also strengthen bounds from indirect detection of particle dark matter, which is largely incompatible with an abundant population of compact objects boosting the annihilation J-factors [94, 95, 96, 97, 98]. Finally, the existence of a PBH population would require specific setups in the early Universe, either deviating from the vanilla slow-roll inflationary scenario [99] or other formation mechanisms based on new physics (see, e.g., Ref. [100]). Another nonprimordial explanation could be subsolar BHs formed from stellar transmutation triggered by dark matter accretion [101, 102, 103, 104, 105, 106, 107], which anyway require new physics.
V.2 NS equation of state
If the SSM objects are identified as light NSs, they would point toward nonstandard formation scenarios, for example, supernovae occurring in sufficiently dense neutron-rich environments [14], since supernova theory struggles to explain subsolar NSs [10].
Furthermore, the large tidal effects can be exploited to constrain the NS EoS. This was anticipated in Fig. 1, showing the projected constraints on $\Lambda_{i}$ and $m_{i}$ inferred from an injection using NS2 with a typical O4/O5 SNR. We selected several widely used representatives from different EoS classes commonly employed to model NS tidal deformability. Specifically, APR [108] and WFF1 [109] are based on microscopic models incorporating realistic nucleon-nucleon interactions, while SLy4 [110] represents mean-field models with effective nuclear interactions, and BSk21 [111] is a phenomenological model fitted to nuclear mass data. For strange quark matter (SQM) stars, we chose SQM3 [41] as a representative model because its tidal deformability results are comparable to those of nonquark families. Notably, SQM1 and SQM2 share the same EoS as SQM3, differing only in the bag constant which only sets a different scale. The models also exhibit varying levels of stiffness: APR and BSk21 are relatively stiff, WFF1 and SLy4 are softer, and SQM3 represents a stiff quark matter EoS.